Parth (parthkohli):

\[\Huge \mathsf{\text{Binary Operations}} \]\[\text{- Lesson to be hosted by nbouscal.} \]

OpenStudy (unklerhaukus):

yes?no?

OpenStudy (anonymous):

Haha alright, why not. Good thing to know.

OpenStudy (anonymous):

Basically, a binary operation is a type of map. Say you have sets A, B, and C, a binary operation maps from \(A\times B\to C\). Most commonly, though, the three sets in question will be the same set. So you'll have for example a map from \(\mathbb{R}\times \mathbb{R}\to \mathbb{R}\).

OpenStudy (anonymous):

The easy examples that you already know are your usual binary operations from arithmetic. Addition, subtraction, multiplication, division, exponentiation.

Parth (parthkohli):

I crave for examples.

OpenStudy (anonymous):

So, with the easy examples you have like 5+6=11, you see you have three elements from R that are mapped via the binary operation of addition. You map (5,6) to {11}. Or if you have 2*4=8, you're mapping (2,4) to {8}.

OpenStudy (anonymous):

But, there are a lot of other binary operations other than those easy examples. For example, composition of functions is a binary operation. If you have f(x) and g(x) then f(g(x)) is a binary operation.

OpenStudy (anonymous):

You would write that one \(f\circ g (x)=f(g(x))\) with \(\circ\) as the operator.

Parth (parthkohli):

So, a binary operation is just something performed on a set to bring it on another? So something like\[7 - 6 = 1 \implies (7,6)\rightarrow \{1 \} \]

OpenStudy (anonymous):

Yeah. The key point is that it's binary, which means it's performed on the product of two sets, and it maps to an element from a single set.

OpenStudy (anonymous):

You can expand it to more than two things, though, by using a certain property that some binary operations have: the associative property.

OpenStudy (anonymous):

Which brings us to an important topic when studying binary operations: properties that they can have. The most important properties are: associative, commutative, identity element, inverses.

Parth (parthkohli):

Example, please?

OpenStudy (anonymous):

Okay we can start with associative property. Associative means that for a binary operation, let's say \(\circ\), you have that \((a\circ b)\circ c=a\circ (b\circ c)\).

Parth (parthkohli):

:O I didn't know that one.

OpenStudy (anonymous):

Easy examples that are associative: addition and multiplication. Easy examples that are not associative: subtraction and division. Also, exponentiation.

Parth (parthkohli):

Yes. Yes. I follow.

Parth (parthkohli):

So is that it?

Parth (parthkohli):

\[\text{Basic operations}\subset \text{Binary operations} \]If I am not wrong.

OpenStudy (anonymous):

After associative you can look at commutative property. Commutative says, for a binary operation \(\circ\), that \(a\circ b=b\circ a\). Again, you can see easy examples with the normal binary operations on scalars. Another good example of noncommutative is matrix multiplication.

OpenStudy (anonymous):

What do you mean by basic operations?

Parth (parthkohli):

Basic operations are \(\times,-,\div, +\)

Parth (parthkohli):

What is \(a \circ b\)?

OpenStudy (anonymous):

Oh, okay. Then yeah, those are a subset of binary operations.

OpenStudy (anonymous):

I'm just using \(\circ\) as a placeholder for an unspecified binary operation.

Parth (parthkohli):

Okay. I see. But \((f\circ g)(x) \ne (g \circ f)(x)\) when they are not inverse functions.

OpenStudy (anonymous):

Right, so composition of functions is not commutative.

OpenStudy (anonymous):

You can also easily see that subtraction and division are not commutative, but addition and multiplication are.

Parth (parthkohli):

Oh, I see. :)

Parth (parthkohli):

Thank you once again! @nbouscal

OpenStudy (anonymous):

There's still more :P

Parth (parthkohli):

:O

Parth (parthkohli):

Go ahead please.

OpenStudy (anonymous):

As I mentioned before, we're also interested in whether binary operations have identity elements, and whether they have inverses.

OpenStudy (anonymous):

An identity element is an element \(e\) such that \(e\circ x = x\circ e = x\) for all x in the set. Can you figure out what the identity elements are for addition and multiplication?

Parth (parthkohli):

Yes. 0 for addition. 1 for multiplication. 1 for division. 0 for subtraction.

OpenStudy (anonymous):

Be careful here. Division and subtraction don't actually have identity elements.

Parth (parthkohli):

Why?\[{x \over 1} = x \]\[x - 0 = x \]

OpenStudy (anonymous):

Notice that I said it had to go both ways. e-x=x-e=x, and e/x=x/e=x. There's no e for either of those that fits that definition.

Parth (parthkohli):

Oh. Yes!

OpenStudy (anonymous):

You'll notice from this definition that the identity element has to commute with every element of the set. That is to say, ex=xe for every x. (We often leave the operation out and just right ex when we mean e\(\circ\)x).

OpenStudy (anonymous):

write*

Parth (parthkohli):

Okay :)

OpenStudy (anonymous):

Should also point out now that division, exponentiation, and some others are sometimes referred to as "partial binary operations" rather than binary operations proper, because they are partial functions. For example, you can't divide by zero.

OpenStudy (anonymous):

So that will explain to some extent why division isn't fitting with any of these properties :P

Parth (parthkohli):

Haha :)

OpenStudy (anonymous):

The last property to talk about for the moment is inverse elements. Basically y is the inverse element of x if xy=yx=e.

OpenStudy (anonymous):

You'll find that some operations have inverse elements for every element of the set, others have inverses for some only, and others have no inverses at all other than the trivial: the identity is always its own inverse. Also you'll notice that to have inverses at all, the operation must have an identity.

OpenStudy (anonymous):

Some other good binary operations to look at are the ones you find in modular arithmetic. That covers it pretty well for now, let me know if you have any other questions.

OpenStudy (anonymous):

(This is good stuff to learn, btw. If you ever take any higher maths at all you'll need to know this for sure.)